HETAV PATEL

Grade 11,

Two identical point mass m move in a circle of radius R due to mutual gravitational attraction.the speed of each mass will be?

Two identical point mass m move in a circle of radius R due to mutual gravitational attraction.the speed of each mass will be?

Grade:12th pass

1 Answers

Vikas TU
3 years ago

Since each particle rotates around the circle of their center of mass , the centripetal acceleration to keep them in the orbit is provided by the mutual gravitational force.

Putting this knowledge into equations,

Centripetal force = Mv^2/R, considering the mass of each particle as 'M' and the rotation speed as 'v'.

Mutual force of gravitational attraction = GM^2/(2R)^2.

Thus, GM^2/4R^2 = Mv^2/R.

We get speed of each particle,

v =(GM/4R)^(1/2).

That's your answer.